# Communicate Less, Synthesize the Rest: Latency-aware Intent-based Generative Semantic Multicasting with Diffusion Models

### v2

@@ 2.87 TFLOPs
Mobile Diffusion-Lite [46]
@@ 38 TFLOPS
In our proposed, generative semantic multicasting scheme, the receivers may either be conventional users equipped with high performance AI processors (e.g. NVIDIA A100 or H100), or mobile users equipped with mobile AI processors (e.g. Apple A18 or M4). On the high performance processors, where we can afford more FLOPs, we use the lightweight diffusion model [18] to maintain the generation quality, while on the mobile processors, we use the recently released MobileDiffusion [46] to enable running the model on-device. In TABLE IV, we report the resulting generation latency $T_{k}^{g}$ for each of these cases. It is concluded that, depending on the deployed model and the device processor, the generation latency is in the range $T_{k}^{g}\in[2,10]$ ms. Hence, we set $T_{1}^{g}=\cdots=T_{K}^{g}=T^{g}$ and plot the ratio of the average per-user latency over the generation latency, i.e. $T^{*}/T^{g}$, for various $T^{g}$ values in this range when $K=10$, in Fig. 9a. This ratio decreases when the generation latency or transmit power budget increases, demonstrating that the generation latency dominates the total latency with increasing $T_{k}^{g}$ and $P_{T}$. Fig. 9b studies the corresponding communication-computation tradeoffs. For example, according to Fig. 9b, a 2 ms increase in the generation latency (6ms >-> 8ms), can be mitigated by increasing the transmit power budget $P_{T}$ (100mw >-> 200mw) to keep the total latency unchanged at $\approx 17$ms.
@@ III-I1 Visual Quality on Cityscapes Street Scenes Dataset
Next, we provide visual quality comparisons with GAN-based synthesis of the non-intended classes in Fig. 18. We replace the diffusion model with a state-of-the-art pretrained GAN model for semantic guided synthesis [47], while all other simulation parameters remain the same. As observed, synthesis with the diffusion model provides photrealistic sharp outputs with better diversity of colors and realism. GAN-based synthesis is blurrier and exhibits noticeable distortion on the class edges, which significantly degrades the overall visual quality.

### v3

@@ 2.87 TFLOPs
Mobile Diffusion-Lite [MobDiff]
@@ 38 TFLOPS
In our proposed, generative semantic multicasting scheme, the receivers may either be conventional users equipped with high performance AI processors (e.g. NVIDIA A100 or H100), or mobile users equipped with mobile AI processors (e.g. Apple A18 or M4). On the high performance processors, where we can afford more FLOPs, we use the lightweight diffusion model [LightDiff] to maintain the generation quality, while on the mobile processors, we use the recently released MobileDiffusion [MobDiff] to enable running the model on-device. In TABLE IV, we report the resulting generation latency $T_{k}^{g}$ for each of these cases. It is concluded that, depending on the deployed model and the device processor, the generation latency is in the range $T_{k}^{g}\in[2,10]$ ms. Hence, we set $T_{1}^{g}=\cdots=T_{K}^{g}=T^{g}$ and plot the ratio of the average per-user latency over the generation latency, i.e. $T^{*}/T^{g}$, for various $T^{g}$ values in this range when $K=10$, in Fig. 9a. This ratio decreases when the generation latency or transmit power budget increases, demonstrating that the generation latency dominates the total latency with increasing $T_{k}^{g}$ and $P_{T}$. Fig. 9b studies the corresponding communication-computation tradeoffs. For example, according to Fig. 9b, a 2 ms increase in the generation latency (6ms >-> 8ms), can be mitigated by increasing the transmit power budget $P_{T}$ (100mw >-> 200mw) to keep the total latency unchanged at $\approx 17$ms.
@@ III-I1 Visual Quality on Cityscapes Street Scenes Dataset
Next, we provide visual quality comparisons with GAN-based synthesis of the non-intended classes in Fig. 18. We replace the diffusion model with a state-of-the-art pretrained GAN model for semantic guided synthesis [oasisgan], while all other simulation parameters remain the same. As observed, synthesis with the diffusion model provides photrealistic sharp outputs with better diversity of colors and realism. GAN-based synthesis is blurrier and exhibits noticeable distortion on the class edges, which significantly degrades the overall visual quality.